OpenStudy (anonymous):

Find the surface area of a sphere with a radius of 5 inches. Leave your answer in terms of pi.

OpenStudy (anonymous):

Im pretty sure its 100 pi

OpenStudy (anonymous):

SA= (4)(pi)(r^2) =4(pi) (5^2) =4(pi)25 =100pi